# A Universal Scheme for Dynamic Partitioned Shortest Path Index

### v2

@@ 3.2. Index Update for Different PSP Strategies
The partitioned shortest path has been widely used, however, how to maintain the index to support its application in dynamic environments has never been discussed. In this section, we propose the partitioned index update algorithms to guarantee the correctness of the partitioned index in dynamic scenarios.
Index Update for Pre-boundary Strategy. As shown in Figure 6-(a), when the weight of edge $e\in G$ changes, we first recalculate the boundary-all pair distance (as the step 1 in index construction) and identify the changed weight $e(b_{i1},b_{i2})$ between boundary vertices in each partition $G_{i}$. Then we need to update the corresponding partition index $L_{i}$ and overlay index $\tilde{L}$.
@@
The indexes of Pre-Boundary Strategy can be correctly updated with the above strategy.
Step 1 takes $O_{u}(nlogn+m)$ time, while the partition and $\tilde{G}$ index takes $max\{O_{u}(G_{i}),O_{u}(\tilde{G})\}$ time to update in parallel, where $O_{u}$ is the update complexity for different indexes.
@@
Index Update for No-boundary Strategy. Since the weight change of inter-edges does not affect the index for each subgraph, we divide index updates into two scenarios as shown in Figure 6-(b). Scenario 1: Inter-edge weight change. When $e\in E_{inter}$ changes, only $\tilde{L}$ needs an update; scenario 2: Intra-edge weight change. When $e\in E_{intra}$ ($e\in E_{j}$) changes, we first update $L_{j}$ and compare the old and new weights of $e(b_{j1},b_{j2})$ between boundary in $G_{j}$. If there is an edge weight update, we need to further update $\tilde{L}$. The update complexity is $max\{O_{u}(G_{i})\}+O_{u}(\tilde{G})$.
@@
The indexes of No-Boundary can be correctly maintained with the above update strategy.
Index Update for Post-boundary Strategy. It is similar to No-Boundary, with an additional judgment and processing as shown in Figure 6-(b). Scenario 1: Intra-edge weight change. Suppose $e\in E_{i}$ changes, we update $G_{i},L_{i}$ and then update $\tilde{G},\tilde{L}$ if any $d_{L_{j}}(b_{j1},b_{j2})$ changes. After that we may need to further update $\{G^{\prime}_{i}\},\{L^{\prime}_{i}\}$ if $d_{G_{i}^{\prime}}(d_{i1},d_{i2})$ and $d_{\tilde{L}}(d_{i1},d_{i2})$ are different; scenario 2: Inter-edge weight change. Suppose $e\in E_{inter}$ changes, we update $\tilde{G},\tilde{L}$ and then update $\{G_{i}^{\prime}\},\{L_{i}^{\prime}\}$. Therefore, the update time complexity is the linear combination of the above procedures: $max\{O_{u}(G_{i})\}+O_{u}(\tilde{G})+O_{q}(\tilde{G})+max\{O_{u}(G_{i})\}$.
@@
The indexes of Post-Boundary Strategy can be correctly maintained with the above update strategy.
Remark 1. To construct the PSP index, Pre-Boundary precomputes the all-pair shortest distance among boundary vertices of each partition, and thus the newly added edges of $G_{i}$ and $\tilde{G}$ (red edges) are of correct edge weight. For example, the edge weight of $e(v_{5},v_{12})$ is 6 which is the path length of the shortest path $p_{G}(v_{5},v_{12})=\left<v_{5},v_{3},v_{10},v_{12}\right>$. While No-Boundary skips the heavy precomputation and only leverage $L_{i}$ to calculate boundary all-pair distance for overlay graph $\tilde{G}$ construction. As a result, the edge weight $w_{\tilde{G}}(v_{5},v_{12})=7$ is incorrect since it only preserves the local distance with corresponding lcoal shortest path $p(v_{5},v_{12})=\left<v_{5},v_{4},v_{12}\right>$ in $G_{3}$. This innovation largely saves the index construction time while drags down the query answering speed. Then Post-Boundary makes up the query efficiency loss by inserting the correct shortest distance value between boundary pairs into each partition (e.g., insert $w(v_{5},v_{12})=6$ to $G^{\prime}_{3}$) and refresh $\{L_{i}\}$ to its correct version $\{L^{\prime}_{i}\}$. In summary, there exists a balance of index performance among the three boundary strategies. In terms of the query processing, the No-Boundary is slower than Pre-Boundary and Post-Boundary, since the boundary vertices and path concatenation should be considered for OD within one partition. In terms of the index construction and update, the No-Boundary and Post-Boundary enjoy faster speed as the boundary all-pair distance computation can be neglected. In terms of the index size, the Post-Boundary needs twice the storage space as that of No-Boundary and Pre-Boundary, even though it has advantages in both query processing and index update. Therefore, no one strategy is better than the others in all aspects and could be selected based on the application scenario. Besides, it is worth noting that we illustrate PSP strategies in the form of edge-cut for clear presentation since the vertex-cut partitions can be easily converted to edge-cut partitions. Due to limited space, the conversion method is presented in extended version (Zhang et al., 2023b).
@@ 3.3. Pruning-based Overlay Graph Optimization
Though it is flexible to select the suitable boundary strategy according to application scenarios, we observe that all boundary strategies are hindered by the extremely dense overlay graph. The density of overlay graph $\tilde{G}$ increases dramatically as the all-pair boundary vertices in each partition are connected during its construction, which badly affects the index performance by slowing down the index construction, query processing, and index update (Zhang et al., 2021a). Then is it possible to improve index performance and decrease the density of $\tilde{G}$ by deleting some unnecessary edges in $\{B_{i}\times B_{i}\}$? As proved in Theorem 1, the shortest distance between boundary vertices is well-preserved in $\tilde{G}$, then could the distance still be preserved in $\tilde{G}$ if some edges are removed? We start exploring this question by categorizing the boundary vertices as followed:
@@
As illustrated in Figure 7-(a), $b_{i3}$ is a half-connected boundary vertex since there exists $v_{i1}\in N(b_{i3}),v_{i1}\notin B_{i}$; and $b_{i1},b_{i2},b_{i4}$ are full-connected boundary vertices. Initially, $\tilde{G}$ can be viewed as composed of multiple complete graphs $\{B_{i}\times B_{i}\}$, which are connected by the inter-edges. Within this dense structure, the global distance between any two boundary vertices $s,t\in B$ is well-preserved. Let us consider a shortest path $p$ with starting vertex $b_{s}\in B_{i}^{F}$ and target vertex $b_{t}\in B$, then the second vertex $v_{0}$ along $p$ is a boundary vertex with $v_{0}\in B,v_{0}\in N(b_{s})$ according to the definition above. Therefore, the global distance information of $p$ will still be kept in $\tilde{G}$ if we delete edges $\{(b_{s},b_{i})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{s})$, since it is impossible for $p$ to pass through these edges. The same principle holds for a shortest path ending with a fully connected boundary vertex $b_{t}\in B_{i}^{F}$ that its distance will also be preserved in $\tilde{G}$ after deleting edges $\{(b_{i},b_{t})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{t})$. Then for the shortest path starting and ending with the half-connected boundary vertices $b_{s},b_{t}\in B_{i}^{H}$, it is necessary to insert an edge $(b_{s},b_{t})$ into $\tilde{G}$ to preserve its distance information, as the second vertex on this path could be a non-boundary vertex which is not contained in the overlay graph. As analyzed above, we can shrink $\tilde{G}$ ($\{B_{i}\times B_{i}\}\cup E_{inter}$) to its slimmer version $\tilde{G}^{\prime}$ (composed of $\{B_{i}^{H}\times B_{i}^{H}\}$, $E_{inter}$ and $B_{i}^{F}$’s adjacent edges). Suppose $b_{i1},b_{i2}\in N(b_{i4})$ and let us only take $G_{i}$ as an example for the clear presentation, the original overlay graph as shown in Figure 7-(b) will be shrunk to a sparser one in Figure 7-(c) after the pruning technique. The following lemma demonstrates that the global distance can still be preserved in the overlay graph with the pruning technique.
@@
$d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$, $\forall(b_{s},b_{t})\in B\times B$.
Therefore, we can enhance the index performance by pruning some edges in $\tilde{G}$ and utilizing its sparser version $\tilde{G}^{\prime}$, while preserving the global distance by referring to Lemma 8.
4. SP-Oriented Partition Method Classification
Having figured out how to achieve efficient query and update in partitioned graphs while guaranteeing the index correctness with the novel PSP strategies, we move on to the remaining two dimensions. As the SP index is well studied and analyzed in Section 2.4. Then, only the mist on the partition method remains: how to choose a partition method to achieve our preferred performance? This is not trivial because: 1) dozens of partition methods with different characteristics were proposed and applied in the past decades; 2) they are classified under different criteria (Buluç et al., 2016), for example, from the perspectives of partition manners (spectral, flow, graph glowing, contraction and multi-level), partition objectives (balance and minimal cut), computation manner (in-memory, distributed and streaming), and cut category (edge-cut and vertex-cut), but it is not clear which criterion is beneficial to SP index; 3) their relationship with (or the influence on) SP index is unknown and has never been studied. Based on these doubts, we propose a novel SP-oriented classification of partition methods as follows:
1) Planar Partition treats partitions equally on one level. It decomposes a graph $G$ into multiple equally-important subgraphs $\{G_{i}|1\leq i\leq k\}(k\in[2,\infty])$ with 1) $\bigcup_{i\in[1,k]}V(G_{i})=V$, $V(G_{i})\cap V(G_{j})=\emptyset$, or 2) $\bigcup_{i\in[1,k]}E(G_{i})=E$, $E(G_{i})\cap E(G_{j})=\emptyset,\forall i,j\in[1,k],i\neq j$. Representative methods include spectral partitioning (Pothen et al., 1990; Barnard and Simon, 1994), graph growing (Diekmann et al., 2000), flow-based partitioning (Sanders and Schulz, 2011; Delling et al., 2011), node-swapping (Kernighan and Lin, 1970; Fiduccia and Mattheyses, 1982), multilevel graph partition (MGP) (Karypis and Kumar, 1998; Barat et al., 2018), etc..
2) Core-Periphery Partition ((Borgatti and Everett, 2000; Rombach et al., 2014; Elliott et al., 2020; Maehara et al., 2014a)) treats the partitions discriminately by taking some important vertices as “Core” and the remaining ones as “Peripheries”. It decomposes a graph $G$ into two distinct parts: a core subgraph $G_{c}$ and $k-1$ peripheral subgraphs $\{G_{i}|1\leq i\leq k-1\}(k\in[2,\infty])$, satisfying $\bigcup_{i\in[1,k-1]}V(G_{i})\cup V(G_{c})=V,V(G_{c})\cap V(G_{i})=\emptyset(\forall 1\leq i\leq k-1)$ and $V(G_{i})\cap V(G_{j})=\emptyset(\forall i,j\in[1,k-1],i\neq j)$. There are two big streams depending on how the core is formed: Core-Tree Decomposition (Maehara et al., 2014b; Li et al., 2020a; Zheng et al., 2022) forms the core through tree decomposition (Robertson and Seymour, 1984), and the resulting periphery part is a set of small-width trees. Specifically, it leverages minimum degree elimination (Xu et al., 2005) to contract vertices of lower degree first such that the graph is contracted from the edge towards the center, generating a set of growing trees (“Peripheries”) around a shrinking while denser graph (“Core”) formed by those non-contracted vertices. The contraction terminates when the width of one tree reaches the previously set threshold; Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022) selects a set of vertices as landmarks, which can be regarded as the core and treats the remaining parts as periphery. It generally works on extremely large graphs where the above tree decomposition is impossible.
3) Hierarchical Partition organizes the network partitions hierarchically and each level is equivalent to planar partitioning. It organizes a graph $G$ hierarchically within $H$ levels where each level $h$ is a planar partitioning $\{G^{h}_{i}\}$ of $G$ and $\exists G^{h-1}_{j}\supset G^{h}_{i},\forall G^{h}_{i}(1\leq h<H)$. Classic methods include HiTi (Jung and Pramanik, 1996), SHARC (Bauer and Delling, 2010), G-Tree (Zhong et al., 2015), G*-Tree (Li et al., 2019a), SMOG (Holzer et al., 2009), and so on.A typical approach to create hierarchical partitions is by utilizing MGP, such as METIS (Karypis and Kumar, 1998) in G-Tree and G*-Tree.
Figure 8. Different SP-Oriented Partitioning
Table 1. SP-Oriented Partition Methods Classification.
Partition Structure Partition Method Cut Category Partition Objectives Spatial-Aware
Category Representative Methods Edge Cut Vertex Cut Balance Minimal Cut
Planar Spectral Partitioning SP (Pothen et al., 1990), RSB (Barnard and Simon, 1994) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Graph Growing Bubble (Diekmann et al., 2000) $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Flow-Based Partitioning PUNCH (Delling et al., 2011) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Geometric Partitioning RCB (Simon, 1991; Huang et al., 1996) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Node-Swapping Heuristic KL (Kernighan and Lin, 1970), FM (Fiduccia and Mattheyses, 1982) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Minimum $k$-cut $k$-cut (Goldschmidt and Hochbaum, 1994) $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
Multilevel Graph Partitioning SCOTCH (Pellegrini and Roman, 1996), METIS (Karypis and Kumar, 1998), KaHyPar (Akhremtsev et al., 2017) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Streaming Edge Partitioning HDRF (Petroni et al., 2015; Mayer and Jacobsen, 2021), CLUGP (Kong et al., 2022) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
In-Memory Edge Partitioning NE (Zhang et al., 2017), HEP (Mayer and Jacobsen, 2021) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
Core-Periphery Core-Tree Decomposition Core-Tree (Maehara et al., 2014b; Li et al., 2020a; Zheng et al., 2022; Elliott et al., 2020) $\checkmark$ $\times$ $\times$ $\times$ $\times$
Sketch Sketch (Potamias et al., 2009; Das Sarma et al., 2010; Gubichev et al., 2010; Tretyakov et al., 2011; Qiao et al., 2012; Qi et al., 2013; Farhan et al., 2018; Wang et al., 2021; Farhan et al., 2022) $\checkmark$ $\times$ $\times$ $\times$ $\times$
Hierarchical Edge-Cut Based HP HiTi (Jung and Pramanik, 1996) $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
SHARC (Bauer and Delling, 2010), G-Tree (Zhong et al., 2015), G*-Tree (Li et al., 2019a) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Vertex-Separator Based HP SMOG (Holzer et al., 2009), ROAD  (Lee et al., 2009, 2010) $\times$ $\checkmark$ $\times$ $\times$ $\times$
Figure 8 shows the partition results of an example graph $G$ with different partitioning methods, where the red vertices represent the corresponding boundary vertices. Specifically, planar partitioning generates four partitions with equivalent vertex size while core-periphery partitioning produces both “Core” and “Periphery” (the periphery of Core-Tree decomposition is a set of small-width trees while Sketch treats the non-core part as periphery. The yellow vertex in each periphery is the “root vertex”). Hierarchical partitioning (HP) organizes partitions hierarchically, and the leaf nodes (the lowest level partition result) may have the same partition result as planar partitioning if the same partitioning method is used. In addition, we summarize and categorize existing partition methods under our SP-oriented classification, as shown in Table 1.
Remark 2. As analyzed in Section 3, the boundary number of each partition will affect the partition method section as it is crucial to the computation complexity of PSP index. Specifically, the planar and hierarchical partitions have no limit to the boundary number, even though reducing the cut size is one of their optimization goals as the path index may suffer from large boundary numbers. Complementary, the core-tree decomposition limits the boundary number to be no larger than the pre-defined bandwidth such that their performance would not be deteriorated by the boundary if we set the bandwidth wisely. Therefore, the planar and hierarchical partitions are better used for small treewidth networks such that balanced partitions and fewer boundaries can be achieved together; while core-periphery is suitable for large treewidth networks as it deliberately limits the boundary vertex number by bandwidth.

### v3

@@ 5.3. Index Update for Different PSP Strategies
The PSP algorithms have been widely used on static networks, however, how to maintain the PSP index to support its application in dynamic environments has never been discussed. In this section, we propose universal index update procedures for different PSP strategies to guarantee the correctness of the partitioned index in dynamic scenarios.
Index Update for Pre-boundary Strategy. As shown in Figure 7-(a), when the weight of edge $e\in G$ changes, we first recalculate the boundary-all pair distance (as the Step 1 in index construction) and identify the changed weight $e(b_{i1},b_{i2})$ between boundary vertices in each partition $G_{i}$. Then we need to update the corresponding partition index $L_{i}$ and overlay index $\tilde{L}$. Step 1 takes $O(|B|\cdot(n\log n+m))$ time, while the partition and $\tilde{G}$ index takes $max\{O_{u}(G_{i}),O_{u}(\tilde{G})\}$ time to update in parallel, where $O_{u}$ is the update complexity for different indexes.
@@
The update procedures of Pre-Boundary Strategy can be correctly deal with any edge weight updates.
Proof.
First of all, we need to recalculate Step 1 to identify the affected edges $e(b_{i1},b_{i2})$ between boundary vertex pairs in each partition. Even though this step is time-consuming, it cannot be skipped since it would be hard to identify the affected edges. For example, suppose the shortest path between the boundary pair $(b_{j1},b_{j2})$ in $G_{j}$ passes through an edge $e\in G_{i}$ with $d_{\tilde{G}}(b_{j1},b_{j2})=d_{0}$. When $e$ increases, we could update $L_{i}$ and then $\tilde{L}$. But $\tilde{L}$ cannot be correctly updated since it could be that $d_{\tilde{G}}(b_{j1},b_{j2})>d_{0}$, such that $d_{\tilde{G}}(b_{j1},b_{j2})$ cannot be refreshed to the correct value. It is because $d_{0}$ contains the old smaller edge weight while cannot be identified, since the shortest distance index always takes the smallest distance value. Then we could select those affected edges $e(b_{i1},b_{i2})$ by comparing their old and new weights. Lastly, we update their corresponding partition index $L_{i}$ and $\tilde{L}$ in parallel. ∎
@@
Index Update for No-Boundary Strategy. Since the weight change of inter-edges does not affect the index for each subgraph, we divide index updates into two scenarios as shown in Figure 7-(b). Scenario 1: Inter-edge weight change. When $e\in E_{inter}$ changes, only $\tilde{L}$ needs an update; Scenario 2: Intra-edge weight change. When $e\in E_{intra}$ ($e\in E_{j}$) changes, we first update $L_{j}$ and compare the old and new weights of $e(b_{j1},b_{j2})$ between boundary in $G_{j}$. If there is an edge weight update, we need to further update $\tilde{L}$. The update complexity is $max\{O_{u}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{u}(\tilde{G})$.
@@
The update procedures of No-Boundary Strategy can be correctly deal with any edge weight updates.
Proof.
In the inter-edge update case, since $e\in\tilde{G},e\notin G_{i},\forall e\in E_{inter}$, the weight change of $e$ could only affect the correctness of $\tilde{L}$. So only $\tilde{L}$ should be checked and updated. In the Intra-edge update case, since $e\in G_{i}$, its weight change will firstly affect $L_{i}$. Then it could affect $\tilde{G}$ as $e_{\tilde{G}}(b_{il_{1}},b_{il_{2}})=d_{L_{i}}(b_{il_{1}},b_{il_{2}})$. So $\tilde{L}$ also needs update if $e_{\tilde{G}}(b_{il_{1}},b_{il_{2}})$ changes after checked. ∎
Index Update for Post-Boundary Strategy. It is similar to No-Boundary with an additional judgment and processing shown in Figure 7-(b). Scenario 1: Intra-edge weight change. Suppose $e\in E_{i}$ changes, we update $G_{i},L_{i}$ and then update $\tilde{G},\tilde{L}$ if any $d_{L_{j}}(b_{j1},b_{j2})$ changes. Then we update $\{G^{\prime}_{i}\},\{L^{\prime}_{i}\}$ if $d_{G_{i}^{\prime}}(d_{i1},d_{i2})$ and $d_{\tilde{L}}(d_{i1},d_{i2})$ are different; Scenario 2: Inter-edge weight change. Suppose $e\in E_{inter}$ changes, we update $\tilde{G},\tilde{L}$ and then update $\{G_{i}^{\prime}\},\{L_{i}^{\prime}\}$. Therefore, the update time complexity is $max\{O_{u}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{u}(\tilde{G})+max\{|B_{i}|^{2}\cdot O_{q}(\tilde{L})\}+max\{O_{u}(G_{i})\}$.
@@
The update procedures of Post-Boundary Strategy can be correctly deal with any edge weight updates.
Proof.
Firstly, we prove the necessity to keep both $\{G_{i}\},\{L_{i}\}$ and $\{G_{i}^{\prime}\},\{L_{i}^{\prime}\}$. Similarly to the Pre-Boundary Strategy, those boundary edges in $\{G_{i}^{\prime}\}$ would keep the old smaller value such that the index could not be correctly updated as explained in Lemma 4. So keeping $\{G_{i}\},\{L_{i}\}$ gives us a chance to update $\tilde{L},\{L_{i}\}$ correctly as proved in Theorem 2. Then, following the No-Boundary Strategy update, we recompute the shortest distance between the all-pair boundaries leveraging $\tilde{L}$ and compare their values on $L_{i}^{\prime}$, then update $G_{i}^{\prime},L_{i}^{\prime}$. ∎
Extension for Structural Updates. We next discuss how to support structural updates (edge/vertex insertion/deletion). For edge deletion, it is equivalent to increasing the edge weight to +$\infty$+ ∞. Vertex deletion can be regarded as deleting all its adjacent edges  (Ouyang et al., 2020; Zhang et al., 2021c, b). Regarding the edge insertion, it is equivalent to decreasing an edge weight from +$\infty$+ ∞ to a certain value (Ouyang et al., 2020; Zhang et al., 2021c, b) and thus we can utilize the edge weight decrease update. In particular, there are three cases for a newly inserted edge $e(u,v)$: 1) $u$ and $v\in G_{i}$ belong to the same partition; 2) $u,v$ belong to different partitions but they are both boundaries; 3) $u,v$ belong to different partitions and at least one of them is not boundary vertex. For case 1), we insert $e$ to $G_{i}$ and leverage edge decrease update to maintain $L_{i}$ (and $\tilde{L}$). For case 2), we update the overlay graph $\tilde{G}$ (and the partition index $L_{i}$ if the endpoints of the new edge belong to the same partition $G_{i}$) and overlay index $\tilde{L}$ by edge weight decrease update. For case 3), supposing $u\notin B$, this case results in a new boundary vertex $u$ and it affects the partition result, so index reconstruction is needed. The vertex insertion is equivalent to a set of edge insertions (Ouyang et al., 2020; Zhang et al., 2021c, b).
Extension for Batch Updates. Our techniques also support batch updates. Since CH and TD initially supports the batch updates, we directly leverage the corresponding techniques (Ouyang et al., 2020; Zhang et al., 2021c) for the overlay and partition index maintenance. While for PLL that only supports streaming updates, a direct approach to process the batch weight update is to leverage the techniques designed for the streaming update. Nevertheless, the all-pair distances computation among boundary vertices could be shared among one batch update for faster index maintenance.
Table 2. Time Complexity of Different PSP Strategies
Opeartions Procedure Complexity
Pre- Boundary Construction $\{B_{i}\times B_{i}\}\rightarrow\{L_{i}\},\tilde{L}$ $O(|B|\cdot(n\log n+m))+max\{O_{c}(G_{i})+O_{c}(\tilde{G})\}$
Query Intra $L_{i}$ $O_{q}(L_{i})$
Inter $L_{i}\oplus\tilde{L}\oplus L_{j}$ $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|\cdot|B_{j}|\cdot O_{q}(\tilde{L}),|B_{j}|\cdot O_{q}(L_{j})\}$
Update $\{B_{i}\times B_{i}\}\rightarrow\{L_{i}\},L$ $O(|B|\cdot(n\log n+m))+max\{O_{u}(G_{i})+O_{u}(\tilde{G})\}$
No- Boundary Construction $\{L_{i}\}\rightarrow\tilde{G}\rightarrow\tilde{L}$ $max\{O_{c}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{c}(\tilde{G})$
Query Intra $L_{i}\oplus\tilde{L}\oplus L_{i}$ $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|^{2}\cdot O_{q}(\tilde{L}))\}$
Inter $L_{i}\oplus\tilde{L}\oplus L_{j}$ $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|\cdot|B_{j}|\cdot O_{q}(\tilde{L}),|B_{j}|\cdot O_{q}(L_{j})\}$
Update $\{L_{i}\}\rightarrow\tilde{G}\rightarrow\tilde{L}$ $max\{O_{u}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{u}(\tilde{G})$
Post- Boundary Construction $\{L_{i}\}\rightarrow\tilde{G}\rightarrow\tilde{L}$$\rightarrow\{B_{i}\times B_{i}\}\rightarrow\{L_{i}\}$ $max\{O_{c}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{c}(\tilde{G})$ $+max\{|B_{i}|^{2}\cdot O_{q}(\tilde{G})\}+max\{O_{u}(G_{i})\}$
Query Intra $L_{i}$ $O_{q}(L_{i})$
Inter $L_{i}\oplus\tilde{L}\oplus L_{j}$ $max\{|B_{i}|\cdot O_{q}(L_{i}),|B_{i}|\cdot|B_{j}|\cdot O_{q}(\tilde{L}),|B_{j}|\cdot O_{q}(L_{j})\}$
Update $\{L_{i}\}\rightarrow\tilde{G}\rightarrow\tilde{L}$$\rightarrow\{B_{i}\times B_{i}\}\rightarrow\{L_{i}\}$ $max\{O_{u}(G_{i})\}+max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\}+O_{u}(\tilde{G})$ $+max\{|B_{i}|^{2}\cdot O_{q}(\tilde{G})\}+max\{O_{u}(G_{i})\}$
Because different SP indexes have different complexities but share the same logical procedure, we use the following notations to represent the logical complexities: $O_{c}$ is the index construction complexity, $O_{q}$ is the query complexity, and $O_{u}$ is the index update complexity. $max$ returns the maximum processing time among all partitions.
We analyze and summarize the time complexity of PSP strategies in terms of index construction, query processing, and index update in Table 2. Notably, the boundary vertex numbers $|B|$ and $|B_{i}|$ are two crucial factors of the index performance. The advantage of the no-boundary and post-boundary strategies over pre-boundary strategy source from the replacement of the time-consuming Dijkstra’s searches ($O(|B|\cdot(m+n\log n)\})$) with index-based all-pair boundary query processing ($O(max\{|B_{i}|^{2}\cdot O_{q}(L_{i})\})$). Therefore, such an optimization generally results in much better performance if the underlying partition index is 2-hop labeling (in which case $O_{q}(L_{i})=w(G_{i})$, where $w(G_{i})$ is the treewidth of $G_{i}$). However, if the query processing is conducted in a search-based manner (e.g., direct search or CH) and the boundary number $|B_{i}|$ is very large, the no-boundary strategy’s performance may dramatically deteriorate. Such a deterioration is even worse for the post-boundary strategy. Nevertheless, no-boundary and post-boundary strategies generally have much better performance than the pre-boundary strategy in most cases.
@@ 5.4. Pruning-based Overlay Graph Optimization
Though it is flexible to select the suitable boundary strategy according to application scenarios, we observe that all boundary strategies are hindered by the extremely dense overlay graph. The density of overlay graph $\tilde{G}$ increases dramatically as the all-pair boundary vertices in each partition are connected during its construction, which badly affects the index performance by slowing down the index construction, query processing, and index update (Zhang et al., 2021a). Then is it possible to improve index performance and decrease the density of $\tilde{G}$ by deleting some unnecessary edges in $\{B_{i}\times B_{i}\}$? We explore this question by categorizing the boundary vertices as follows:
@@
As illustrated in Figure 8-(a), suppose partition $G_{i}$ has five boundary vertices, where $b_{1}$, $b_{2}$, and $b_{4}$ are full-connected boundary vertices since their neighbors are all boundary vertices while $b_{3}$ and $b_{5}$ are half-connected boundary vertex since they have neighbors that are non-boundary vertices ($v_{1}$ and $v_{2}$). We observe that the shortest paths from full-connected boundary vertices to other boundary vertices always pass through at least one of their boundary neighbors. Let us consider the shortest path $p$ from $b_{4}\in B_{i}^{F}$ to any other boundary vertex $b\in B$, then the second vertex $v$ along $p$ must be a boundary vertex with $v\in N(b_{4})$. Therefore, the global distance information of $p$ will still be kept in $\tilde{G}$ if we delete edges $\{(b_{4},b_{i})\}$ with $b_{i}\in B_{i},b_{i}\notin N(b_{4})$, since it is impossible for $p$ to pass through these edges. By contrast, for the shortest path starting and ending with the half-connected boundary vertices $b_{s},b_{t}\in B_{i}^{H}$, it is necessary to insert an edge $(b_{s},b_{t})$ into $\tilde{G}$ to preserve its distance information, as the second vertex on this path could be a non-boundary vertex which is not contained in the overlay graph. As analyzed above, we can shrink $\tilde{G}$ ($\{B_{i}\times B_{i}\}\cup E_{inter}$) to its slimmer version $\tilde{G}^{\prime}$ (composed of $\{B_{i}^{H}\times B_{i}^{H}\}$, $E_{inter}$ and $B_{i}^{F}$’s adjacent edges). Let us only take $G_{i}$ as an example for the clear presentation, the original overlay graph as shown in Figure 8-(b) will be shrunk to a sparser one in Figure 8-(c) after the pruning technique, reducing 6 redundant shortcuts. In the following Lemma, we prove the global distance can still be preserved in the overlay graph after pruning.
@@
$\forall b_{s},b_{t}\in B,d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$.
Proof.
We denote the concise form of $p_{G}(b_{s},b_{t})$ by only taking the borders as $\{b_{s}=b_{0},\dots,b_{t}=b_{h}\},(0<j<h)$.
When $h=1$, if either $b_{s}\in B^{F}$ or $b_{t}\in B^{F}$, then $d_{G}(b_{s},b_{t})=w_{G}(b_{s},b_{t})$ holds with $b_{s}\in N(b_{t})$ by referring to definition 7. Since $(b_{s},b_{t})\in\tilde{G}^{\prime}$, $d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$ holds. The same applies when $b_{s},b_{t}\in B^{H}$ with two endpoints in different subgraphs $b_{s}\in B_{i},b_{t}\in B_{j}(i\neq j)$. If $b_{s},b_{t}\in B^{H}$ with $b_{s},b_{t}\in B_{i}$, edge $(b_{s},b_{t})$ are inserted into $\tilde{G}^{\prime}$ with their global shortest distance in both Pre-Boundary and Post-Boundary strategies, so $d_{G}(b_{s},b_{t})=d_{\tilde{G}^{\prime}}(b_{s},b_{t})$ naturally holds. In the No-Boundary strategy, though only the local shortest distance is inserted in the overlay graph with $d_{G_{i}}(b_{s},b_{t})=d_{\tilde{G}^{\prime}}(b_{s},b_{t})$, it holds that $d_{G}(b_{s},b_{t})=d_{G_{i}}(b_{s},b_{t})$ since there exists no other boundary vertex besides two endpoints in $p_{G}(b_{s},b_{t})$ and it only pass through $G_{i}$, so $d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$ holds.
When $h>1$, the shortest path distance is accumulated as $d_{G}(b_{s},b_{t})=\sum_{j=0}^{h-1}d_{G}(b_{j},b_{j+1})$. Since $d_{G}(b_{j},b_{j+1})=d_{\tilde{G}^{\prime}}(b_{j},b_{j+1})$ holds for $0\leq j<h$ by referring to the case when $h$=1, we can get that $d_{G}(b_{s},b_{t})=d_{\tilde{G}^{\prime}}(b_{s},b_{t})$. So we prove that $d_{G}(b_{s},b_{t})=d_{\tilde{G^{\prime}}}(b_{s},b_{t})$ holds for all scenarios and boundary strategies. ∎
Therefore, we enhance the index performance by utilizing the sparser $\tilde{G}^{\prime}$. Note that Lemma 8 also holds with edge weight updates since this optimization only depends on the structure of $\tilde{G}$.